OpenStudy (anonymous):

3(x + 1) = −2(x − 1) − 4

OpenStudy (theviper):

\[\color{red}{3x+3=-2x+2-4}\]

OpenStudy (theviper):

\[\color{green}{3x+3=-2x-2}\]

OpenStudy (theviper):

\[\color{orange}{3x+2x=-3-2=-5}\]

OpenStudy (theviper):

\[\color{green}{5x=-5}\]

OpenStudy (theviper):

\[\LARGE{x=\frac{-5}{5}=-1}\]

OpenStudy (anonymous):

lol What's with the different colors? 3(x + 1) = −2(x − 1) − 4 Distribute: 3x + 3 = -2x + 2 - 4 Add like terms: 3x + 3 = -2x - 2 Bring like terms to separate sides: 5x = -5 x = -5/5 x = -1

OpenStudy (theviper):

I wrote it colorfully as the asker can easily understand:)